OpenStudy (anonymous):

A cube of silver is 5.00 cm on a side and has a mass of 1312.5 g. What is the density of silver? Please help explain how to do it, don't give the answer!!

OpenStudy (aaronq):

density is equal to a mass in a certain volume density = mass/volume

OpenStudy (anonymous):

So 1312.5g/5.00cm?

OpenStudy (anonymous):

You have to find the volume of the cube first. \[S^3 = 5.00^3 = 125cm^3\] Then you can use the density formula